# Revisiting XL-MIMO Channel Estimation: When Dual-Wideband Effects Meet Near Field

### v3

@@ VI Simulation Results
Notations Parameters
Number of BS antenna $N_{\mathrm{R}}$ 256
Number of RF chain $N_{\mathrm{RF}}$ 16
Carrier frequency $f_{c}$ 30GHz
Number of pilot carriers $K$ 64
System bandwidth $f_{s}$ 1.6GHz
Number of channel path $L$ 4
Angle of arrival $\vartheta_{l}$ $\mathcal{U}(-\pi/2,\pi/2)$
Distance between BS and UE or scatterers $r_{l}$ [5, 50]m
Proportion of visible antenna elements $\rho_{l}$ (0,1]
@@
  • UAMP-SBL[40]: An improved version of the StdSBL algorithm that leverages the unitary AMP framework to perform the E-step and variational message passing to achieve the updates of prior parameters.
@@ VI-B NMSE versus SNR and Pilot Symbol Number
Fig. 6 illustrates the NMSE performance of various algorithms as a function of SNR for $P=8$. The results reveal that algorithms such as SOMP [36], StdSBL [46], and UAMP-SBL [40], which fail to account for both global block sparsity and local common-delay sparsity, exhibit significantly poorer NMSE performance compared to algorithms like PC-SBL, VSP, and the proposed MMV-HMP. Among block-sparsity prior-based methods, the PC-SBL algorithm leverages a pattern-coupled Gaussian prior model to effectively capture local common-delay sparsity. However, it does not consider global block sparsity. In contrast, the VSP algorithm employs a MRF-based prior to model global block sparsity but lacks the ability to incorporate local common-delay sparsity. In contrast, the proposed MMV-HMP algorithm overcomes these limitations by utilizing a tailored column-wise hierarchical prior that simultaneously incorporates both MRF structure and precision parameter sharing mechanism. This comprehensive modeling approach ensures consistently superior NMSE performance across the entire SNR range under consideration.
@@ VI-C NMSE versus Path Number
Fig. 8 illustrates the NMSE performance of various algorithms as a function of the number of propagation paths, with $P=8$ and $\mathrm{SNR}=2$dB. The SOMP algorithm exhibits significant sensitivity to prior knowledge of the number of paths, with its performance degrading sharply when the assumed and actual path numbers differ. In contrast, Bayesian inference-based methods, leveraging sparsity-promoting prior models, adapt effectively to variations in the number of paths, maintaining consistent and robust performance across all scenarios. Notably, the proposed MMV-HMP algorithm consistently outperforms other Bayesian methods, demonstrating superior capability in handling dynamic sparse scattering environments.
@@ VII Conclusions
In this paper, we have addressed the channel estimation problem in XL-MIMO systems, taking into account the spherical wavefront effects, SnS properties, and dual-wideband effects. We began by rigorously quantifying the angular and delay spread properties of SnS dual-wideband channels in the angular-delay domain and revealed their inherent global block sparsity and local common-delay sparsity. To exploit this structured sparsity, we proposed a computationally efficient MMV-HMP algorithm. The algorithm performs message updates along the factor graph by selectively applying either the SPA or VMP rules, depending on the structure and tractability of each factor node. Simulation results demonstrate that the proposed MMV-HMP algorithm consistently outperforms existing baseline methods, achieving robust estimation performance with moderate computational complexity. Furthermore, an ablation study confirms the effectiveness of the proposed column-wise hierarchical prior, validating its contribution to improved estimation accuracy.
@@ Appendix A Proof of Lemma 1
The proof of Lemma 1 is divided into two parts. First, we demonstrate that the angular-frequency representation $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$ exhibits row-wise sparsity. Next, we show that the spatial-delay representation $\boldsymbol{\Theta}(\varphi_{,}\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ exhibits column-wise sparsity. Without loss of generality, we assume that $\cos\vartheta>0$.
We begin by examining the angular transformation of $\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$. For a fixed $k$, the angular transformation of $\boldsymbol{\theta}(f_{k},\varphi_{l},\psi_{l})$ is given by
$[\boldsymbol{\Xi}_{l,\mathrm{A}}]_{m,k}=\sum_{n=1}^{N}\mathrm{e}^{\mathrm{j}\frac{2\pi}{N}mn}\mathrm{e}^{\mathrm{j}2\pi\left(a_{l,k}n-\frac{b_{l,k}}{2}n\right)n}.$ (A.1)
According to (21) and (22), it can be obtained that $[\boldsymbol{\Xi}_{l,\mathrm{A}}]_{:,k}$ demonstrates block sparsity with its prominent spatial frequency components concentrated from $i_{l,e}(k)$ to $i_{l,s}(k)$. Extending this observation across all $K$ subcarriers, the overall spatial frequency range of significant components spans from $i_{l,e}^{\mathrm{min}}=\min_{k}i_{l,e}(k)$ to $i_{l,s}^{\mathrm{max}}=\max_{k}i_{l,s}(k)$. As a result, $\boldsymbol{\Xi}_{\mathrm{A}}$ is characterized as a row-wise block matrix
@@
$\lim_{K\to\infty}\left|[\boldsymbol{\Xi}_{l,n}]_{n,m}\right|\propto\sqrt{K}\delta\left(t_{l,n}-\frac{m}{K}\right).$ (A.4)
From (A.4), when $K\to\infty$, $[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,:}$ can be approximated as a delta function centered at $m=t_{l,n}K$. However, due to the finite sampling size $K$ in practice, we consider the range from $J_{l,e}(n)=(t_{n}-{2}/{K}+1/2)K$ to $J_{l,s}(n)=(t_{n}+{2}/{K}+1/2)K$ to collect the delay indices near the peak effectively. Extending this observation across all $N$ antennas, $\boldsymbol{\Xi}_{\mathrm{D}}$ is characterized as a column-wise block matrix
$\boldsymbol{\Xi}_{l,\mathrm{D}}=\left[\mathbf{0}_{N\times(J_{l,e}-1)},\mathbf{L}_{N\times(J_{l,s}-J_{l,e}+1)},\mathbf{0}_{N\times(K-J_{l,s})}\right],$ (A.5)
where $J_{l,s}=\min_{n}J_{l,s}(n)$ to $J_{l,e}=\max_{n}J_{l,e}(n)$ with $n\in\boldsymbol{\phi}$, and $\mathbf{L}$ is a non-zero matrix.
@@
Moreover, based on the definition $\boldsymbol{\Xi}_{l}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$, it follows that, for a fixed column index $k$, $\boldsymbol{\Xi}_{l}$ characterizes the angular distribution corresponding to a specific propagation delay, which is determined by the spatial frequencies across different subcarriers. Accordingly, within the region $\mathcal{A}_{l}$, each column captures the angular response for a given delay, and the width of significant spatial frequency components is approximately determined by the system width and the start spatial frequency, i.e., $i_{l,k}N{f_{s}}/{f_{c}}<I_{l,s}-I_{l,e}$, where $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$ denotes the starting spatial frequency of the $k$-th column. As a result, the nonzero components of $\boldsymbol{\Xi}_{l}$ are confined within a narrow band in each column, leading to a distinct column-wise clustered sparsity pattern.
@@ Appendix B A Proof of Theorem 1
$\mathbf{H}_{l,p}=\alpha_{l}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p})\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (B.1)
@@
$\mathcal{B}_{l,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l,1}+N\epsilon_{l,p},N),\right.$ (B.2)
@@
Here, $\mathcal{A}_{l}$ is the non-zero region for $\boldsymbol{\Xi}_{l}$ given by (25), and mod$(a,m)$ is the modulus of $a$ for $m$. Considering all $P$ frequency components, the significant entries of $\mathbf{X}_{l}$ would be localized within the region $\mathcal{B}_{l}\triangleq\cup_{p}\mathcal{B}_{l,p}$. Thus, the proof is completed.

### v4

@@ VI Simulation Results
Notations
Parameters
Number of BS antenna $N_{\mathrm{R}}$
256
Number of RF chain $N_{\mathrm{RF}}$
16
Carrier frequency $f_{c}$
30GHz
Number of pilot carriers $K$
64
System bandwidth $f_{s}$
1.6GHz
Number of channel path $L$
4
Angle of arrival $\vartheta_{l}$
$\mathcal{U}(-\pi/2,\pi/2)$
Distance between BS and UE or scatterers $r_{l}$
[5, 50]m
Proportion of visible antenna elements $\rho_{l}$
(0,1]
@@
  • UAMP-SBL[40]: An improved version of the StdSBL algorithm that leverages the UAMP framework to perform the E-step and variational message passing to achieve the updates of prior parameters.
@@ VI-B NMSE versus SNR and Pilot Symbol Number
Fig. 6 illustrates the NMSE performance of various algorithms as a function of SNR for $P=8$. The results reveal that algorithms such as SOMP [36], StdSBL [46], and UAMP-SBL [40], which fail to account for both global block sparsity and local common-delay sparsity, exhibit significantly poorer NMSE performance compared to algorithms like PC-SBL, VSP, and the proposed MMV-HMP. Among block-sparsity prior-based methods, the PC-SBL algorithm leverages a pattern-coupled Gaussian prior model to effectively capture local common-delay sparsity. However, it does not consider global block sparsity. In contrast, the VSP algorithm employs a MRF-based prior to model global block sparsity but lacks the ability to incorporate local common-delay sparsity. Overall, the performance degradation of these SBL-based baselines stems from the mismatch between their prior models and the structured sparsity of XL-MIMO channels. In contrast, the proposed MMV-HMP algorithm overcomes these limitations by utilizing a tailored column-wise hierarchical prior that simultaneously incorporates both MRF structure and precision sharing mechanism. This comprehensive modeling approach ensures consistently superior NMSE performance across the entire SNR range under consideration.
@@ VI-C NMSE versus Path Number
Fig. 8 illustrates the NMSE performance of various algorithms as a function of the number of propagation paths, with $P=8$ and $\mathrm{SNR}=2$dB. The SOMP algorithm exhibits significant sensitivity to prior knowledge of the number of paths, with its performance degrading sharply when the assumed and actual path numbers differ. In contrast, Bayesian inference-based methods, leveraging sparsity-promoting prior models, adapt effectively to variations in the number of paths, maintaining consistent and robust performance across all scenarios. Notably, the proposed MMV-HMP algorithm consistently outperforms other Bayesian methods, demonstrating superior capability in handling dynamic scattering environments.
@@ VII Conclusions
In this paper, we have addressed the channel estimation problem in XL-MIMO systems, taking into account the spherical wavefront effects, SnS properties, and dual-wideband effects. We began by rigorously quantifying the angular and delay spread properties of SnS dual-wideband channels in the angular-delay domain and revealed their inherent global block sparsity and local common-delay sparsity. To exploit this structured sparsity, we proposed a computationally efficient MMV-HMP algorithm. Simulation results demonstrate the superiority of the MMV-HMP algorithm in both computational complexity and estimation performance. Furthermore, an ablation study confirms the effectiveness of the proposed column-wise hierarchical prior, validating its key contribution to enhanced estimation accuracy.
@@ Appendix A Proof of Lemma 1
The proof of Lemma 1 is divided into two parts. First, we demonstrate that the angular-frequency representation $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$ exhibits row-wise sparsity. Next, we show that the spatial-delay representation $\boldsymbol{\Theta}(\varphi_{,}\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ exhibits column-wise sparsity.
For a fixed $k$, the angular transformation of $\boldsymbol{\theta}(f_{k},\varphi_{l},\psi_{l})$ is given by
$\boldsymbol{\Xi}_{l,\mathrm{A}}]_{m,k}=\sum_{n=1}^{N}\mathrm{e}^{\mathrm{j}\frac{2\pi}{N}mn}\mathrm{e}^{\mathrm{j}2\pi\left(a_{l,k}n-\frac{b_{l,k}}{2}n\right)n}.$ (A.1)
According to (21) and (22), it can be obtained that $\boldsymbol{\Xi}_{l,\mathrm{A}}]_{:,k}$ demonstrates block sparsity with its prominent spatial frequency components concentrated from $i_{l,e}(k)$ to $i_{l,s}(k)$. Extending this observation across all $K$ subcarriers, the overall spatial frequency range of significant components spans from $i_{l,e}^{\mathrm{min}}=\min_{k}i_{l,e}(k)$ to $i_{l,s}^{\mathrm{max}}=\max_{k}i_{l,s}(k)$. As a result, $\boldsymbol{\Xi}_{\mathrm{A}}$ is characterized as a row-wise block matrix
@@
$\lim_{K\to\infty}\left|[\boldsymbol{\Xi}_{l,D}]_{n,m}\right|\propto\sqrt{K}\delta\left(t_{l,n}-\frac{m}{K}\right).$ (A.4)
From (A.4), when $K\to\infty$, $\boldsymbol{\Xi}_{l,D}]_{n,:}$ can be approximated as a delta function centered at $m=t_{l,n}K$. However, due to the finite sampling size $K$ in practice, we consider the range from $j_{l,s}(n)=(t_{n}-{2}/{K}+1/2)K$ to $j_{l,e}(n)=(t_{n}+{2}/{K}+1/2)K$ to collect the delay indices near the peak effectively. Extending this observation across all $N$ antennas, $\boldsymbol{\Xi}_{\mathrm{D}}$ is characterized as a column-wise block matrix
$\boldsymbol{\Xi}_{l,\mathrm{D}}=\left[\mathbf{0}_{N\times(J_{l,s}-1)},\mathbf{L}_{N\times(J_{l,e}-J_{l,s}+1)},\mathbf{0}_{N\times(K-J_{l,e})}\right],$ (A.5)
where $J_{l,s}=\min_{n}j_{l,s}(n)$ and $J_{l,e}=\max_{n}j_{l,e}(n)$ with $n\in\boldsymbol{\phi}_{l}$, and $\mathbf{L}$ is a non-zero matrix.
@@
Moreover, for a fixed column index $k$, $\boldsymbol{\Xi}_{l}$ characterizes the angular distribution corresponding to a specific propagation delay, which is determined by the spatial frequencies across different subcarriers. Accordingly, within the region $\mathcal{A}_{l}$, each column captures the angular response for a given delay, and the width of significant spatial frequency components is approximately determined by the system bandwidth and the start spatial frequency, i.e., $i_{l,k}N{f_{s}}/{f_{c}}$, where $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$ denotes the starting spatial frequency of the $k$-th column. As a result, the nonzero components of $\boldsymbol{\Xi}_{l}$ are confined within a narrow band in each column, leading to a distinct column-wise clustered sparsity pattern.
@@ Appendix B A Proof of Theorem 1
$\mathbf{H}_{l,p}=\alpha_{l}c_{l,p}\left(\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p})\mathbf{a}^{\mathrm{T}}(\tau_{l})\right)\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (B.1)
@@
$\mathcal{B}_{l,p}$ $=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l,1}+N\epsilon_{l,p},N),\right$. (B.2)
@@
Here, $\mathcal{A}_{l}$ is the non-zero region for $\boldsymbol{\Xi}_{l}$ given by (25), and mod(a,m)(a,m) is the modulus of $a$ for $m$. Considering all $P$ frequency components, the significant entries of $\mathbf{X}_{l}$ would be localized within the region $\mathcal{B}_{l}\triangleq\cup_{p}\mathcal{B}_{l,p}$. Thus, the proof is completed.